OpenStudy (anonymous):

What are the x and y intercepts of the quadratic equation y= (x-3)^2 -25? 1) x: (-2,0) and (8,0) y: (0,-16) 2) x: (2,0) and (-8,0) y: (0,-16) 3) x: (-2,0) and (8,0) y: (0,-25) 4) x: (2,0) and (-8,0) y: (0,-25)

OpenStudy (anonymous):

y intercept: put x = 0 and find y you get \[y=(0-3)^2-25=9-25=-16\]

OpenStudy (anonymous):

x=8,-2, y=-16

OpenStudy (anonymous):

to find the x - intercept set y = 0 solve for x you get \[(x-3)^2-25=0\] \[(x-3)^2=25\] \[x-3=5\] or \[x-3=-5\] \[x=8\] or \[x=-2\] so x - intercepts are 5 and -2

OpenStudy (anonymous):

sorry i mean of course 8 and -2